OpenStudy (anonymous):

An eastbound train is traveling 25.5m/s when the engineer applies the brake for 17.0s to reduce the speed of the train to 16.4m/s what is the average accelerating during the 17.0s period?

OpenStudy (anonymous):

by the equation (v-u)/t=a we get a=(25.5-16.4)/17=0.535 and as acceleration is negative because the train will retard due to the application of brakes, the acceleration=-0.535